OpenStudy (anonymous):

Which fractions are equivalent to -1/5 ? Choose all answers that are correct. A. -10/-25 B. -2/10 <-(In that only the 2 is negative) C. -4/-25 D.5/ -25

OpenStudy (jojokiw3):

But multiplying both the numerator and the denominator with the same value, the value of the number itself is not changed. Thus for example. \[\frac{ 1 }{ 10 } = \frac{ 2 }{ 20 }\] How I got that was by multiplying \[\frac{ 1 }{ 10 }(\frac{ 2 }{ 2 })\] which equals \[\frac{ 2 }{ 20 }\] Now just simply remember that multiplying a - with + or a + with - equals a -. Then test all the possible answers.

OpenStudy (anonymous):

So its C. & B.?